OpenStudy (buttermequeasy):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

OpenStudy (buttermequeasy):

HELP!

OpenStudy (mysesshou):

I think someone did this already ? slope intercept form is y = m*x + b slope = m = -6 y intercept = b = 9 plug into equation y = -6*x + 9 and that is it. :)